OpenStudy (anonymous):

The number of milligrams of a drug that remains in a patient’s system after t hours is given by the function A(t) = Ie^rt. Juan was given 500 milligrams of medicine which leaves his bloodstream at a rate of 25%. How much of the medicine remains in his system after 6 hours? 150.62 mg 133.90 mg 111.57 mg 99.20 mg

OpenStudy (anonymous):

@kropot72 could u help again

OpenStudy (anonymous):

@lordcyborg help

OpenStudy (anonymous):

@blurbendy help

OpenStudy (anonymous):

Is that 'L' ????

OpenStudy (anonymous):

I mean A(t) = L e^rt ????

OpenStudy (anonymous):

@lordcyborg

OpenStudy (anonymous):

@Jim766 need help

OpenStudy (anonymous):

Its i which is 500

OpenStudy (anonymous):

not L

OpenStudy (anonymous):

hello??

OpenStudy (anonymous):

ya ya sorry

OpenStudy (anonymous):

See u get l=500 mg r = 1/4 and t= 6*3600 (in seconds) So just put the value and get the answer...... Give it a try

OpenStudy (anonymous):

please just need the answer and no time to put it into a calculator

OpenStudy (anonymous):

y man plzz try..........thats not the way cmon

OpenStudy (anonymous):

i just need the answer

OpenStudy (anonymous):

not the formula

OpenStudy (anonymous):

I cant help this way man........This website is made for helping and teaching the correct methods..... not to just give away answers.... I've been warned by mods twice for that

OpenStudy (anonymous):

plzz

OpenStudy (anonymous):

Jim will u give me the answer

OpenStudy (anonymous):

i always give out answers

OpenStudy (anonymous):

at leaset give me the formula with the numbers plugged in

OpenStudy (anonymous):

ok i'm trying By the way : A=(500)*e^(3/2)

OpenStudy (anonymous):

so the answer is...

OpenStudy (anonymous):

wts e^(3/2)

OpenStudy (anonymous):

@Jim766 will u help or just sit there

OpenStudy (anonymous):

4.48

OpenStudy (anonymous):

right?

OpenStudy (jim766):

sorry had to look it up...formula should be to the ^-rt 500 * e ^-(.25*6)

OpenStudy (anonymous):

Hey @pvs285 it should be e^(-rt)

OpenStudy (anonymous):

ya @Jim766

OpenStudy (anonymous):

yup

OpenStudy (jim766):

111.57

OpenStudy (anonymous):

so the answer is...

OpenStudy (anonymous):

thank you

OpenStudy (jim766):

yw

OpenStudy (anonymous):

what does that mean by the way

OpenStudy (anonymous):

yw : you r welcome

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

god what are u typing

OpenStudy (jim766):

if you take a drug, using this formula, you can tell much of it is metabolized for a certain amount of time, and how much is still in your system in this case 6 hours later 111.57 mg of the medicine is still in your system...and working I guess...can tell you you need to take another dose...can make sure you dont take to much..

OpenStudy (jim766):

sorry for the book..

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

bey